OpenStudy (anonymous):

Find the vertex of the quadratic equation: y = 2x2 + 12x − 4. (−3, 50) (3, 50) (−3, −22) (−3, −58)

OpenStudy (anonymous):

What's a vertex?

OpenStudy (anonymous):

Hint: derivative

OpenStudy (campbell_st):

the vertex is where the curve changes from decreasing to increasing... or vice versa... easist way is \[x = -\frac{b}{2a}\] in your equation you have a = 2 and b = 12 so \[x = \frac{-12}{4}\] simplify it to get the x value of the vertex. then substitute it into the original equation to find the y value

OpenStudy (anonymous):

I'm sorry....are you still confused?

OpenStudy (anonymous):

lol sorry but yes.

OpenStudy (anonymous):

what will x be ?

OpenStudy (anonymous):

umm 3

OpenStudy (anonymous):

can you tell me what \[\frac{-12}{4}\] is?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

great now that we know that \[x=-3\] We want to know what y is \[y = 2x^2 + 12x − 4\] All you have to do is plug in \[x=-3\] to the top equation and solve for y After doing that you will get your (x,y)

OpenStudy (anonymous):

soo y= -22...?

OpenStudy (anonymous):

yes so it's (-3,-22)